# Combinatorial Selection with Costly Information

### v1

@@ Abstract
We develop a framework for approximate optimization of bandit superprocesses that applies to arbitrary processes with a matroid (and in some cases, more general) feasibility constraint. Our framework establishes a bound on the optimal cost through a novel cost amortization; it then couples this bound with a notion of local approximation that allows approximate solutions for each component MDP in the superprocess to be composed without loss into a global approximation.
We use this framework to obtain approximately optimal solutions for several variants of bandit superprocesses for both maximization and minimization. We obtain new approximations for combinatorial versions of the previously studied Pandora’s Box with Optional Inspection and Pandora’s Box with Partial Inspection; as well as approximation algorithms for a new problem that we call the Weighing Scale problem.
@@ 1 Introduction
Consider a biotechnology company investing in multiple drug discovery projects. Each individual project can involve a multitude of decisions with uncertain outcomes. Typically very few projects will make it to trials and FDA approval at which point they can begin to yield rewards. The company needs to determine how to distribute resources across these projects without knowing upfront which one will succeed and to what extent. Many decision making problems call for upfront investment in information acquisition (such as how much reward a project will generate) before optimizing over multiple options (such as which project to seek FDA approval for). Some other examples include an oil company’s search for a good drilling site; a construction company exploring different design ideas for a project; a manufacturer performing market research before determining which products to make and in what quantity, etc.
As a toy example of costly information acquisition in the context of optimization, consider what we call the Weighing Scale problem. The algorithm is presented with $n$ alternatives with unknown costs, $X_{1},\cdots,X_{n}$, of which it needs to pick a cheap one. Each cost $X_{i}$ is distributed independently according to a known distribution. The only way the algorithm can learn any further information about the $X_{i}$’s is to use a weighing scale at the additional cost of $1$ per use: for any index $i$ and target $t$, the weighing scale returns $\mathbbm{1}[X_{i}\leq t]$. What information, if any, should the algorithm gather before making its selection so as to minimize the expected cost of the chosen alternative plus the number of weighings?
In this paper we study a class of problems that we call Costly Information Combinatorial Selection (henceforth, CICS), of which Weighing Scale is an example. Our goal is to pick a subset of $n$ alternatives with stochastic values satisfying a feasibility constraint. We can obtain information about the value of any individual alternative through a sequence of costly actions tailored to that alternative. In the minimization version of the problem our objective is to minimize the total value of the chosen alternatives plus the cost of information acquisition. In the maximization version we maximize the total value of the chosen subset minus the cost of information acquisition.
The algorithmic challenge of CICS stems from two levels of decision-making that influence each other – which alternative to explore in each step; as well as how to explore the chosen alternative (e.g. which target values $t$ to use in the Weighing Scale setting). A policy can, in general, make these choices adaptively based on the information obtained on previous steps.
A classic example of the CICS is Weitzman’s Pandora’s Box problem. Here each alternative is hidden inside a closed box; the goal is to select a single alternative; the algorithm can observe the value of the alternative by opening, a.k.a. “inspecting”, each box at a cost; and any alternative must be inspected before being selected. Pandora’s Box admits an astonishingly simple optimal algorithm: each box is assigned an index based on its cost and value distribution; for minimization, boxes are inspected in increasing [^4] For the maximization version of the problem, boxes are opened in decreasing order of index. order of index until the algorithm finds a value that is smaller than all remaining indices and accepts that value. The indices do not depend on the algorithm’s stochastic trajectory; the order of information acquisition is therefore non-adaptive. Essentially, the indices capture all of the salient information about each individual box that can then be used by the algorithm to determine the order in which to explore the boxes.
Interestingly, this optimality structure extends to a large class of feasibility constraints over Pandora’s Boxes: Singla (2017) showed that Weitzman’s indices and the corresponding greedy ordering algorithm achieve approximate optimality for essentially all selection problems that admit greedy-style (a.k.a. frugal) approximation algorithms.
However, the landscape changes if the algorithm has multiple modes of information acquisition available for each alternative. For example, if inspection becomes optional, that is, if the algorithm is allowed to select an alternative without observing its value, the optimal algorithm is no longer non-adaptive and the problem becomes NP-hard. Much effort has been devoted recently towards developing approximations for variants of Pandora’s Box, including the optional inspection version and the so-called partial inspection version. See, for example, the survey by Beyhaghi and Cai (2024). These results are generally tailored to the specifics of the variants they study and do not extend to other models of information acquisition. Moreover, they do not always generalize to combinatorial settings.
A natural question is whether index-based policies can be extended to other CICS settings: can we distill the essence of the information acquisition process for each alternative into a Weitzman-style index, and build an approximation algorithm based on it? This is the question we seek to answer in this paper.
We develop a framework for designing approximately-optimal policies for both the minimization and maximization versions of CICS. Our framework has two components. First we develop a bound on the cost of the optimal solution based on a novel cost amortization technique that generalizes the Gittins index to Markov Decision Processes. This bound allows us to disentagle the exploration of different alternatives into individual “local” problems. We then show that a certain notion of local approximation applied to each individual exploration problem can be combined via an index-based policy to obtain a global approximation guarantee. This approach is a strengthening and generalization of the “Local Hedging” technique developed by Scully and Doval (2024) in the context of Pandora’s Box with optional inspection.
We instantiate our framework with new approximation results for previously studied variants of Pandora’s Box as well as the new weighing scale problem described above. Our framework and results extend seamlessly to combinatorial selection problems over matroid constraints as well as to combinations of different variants of information acquisition. For example, when given a selection problem with some weighing-scale-alternatives and some optional-inspection-alternatives, our approach can effortlessly combine local approximations for each variant to obtain a combined guarantee for the whole problem.
Bandit superprocesses and our cost amortization framework
Selection problems with costly information are closely related to a class of sequential decision-making problems called bandit superprocesses, which generalize Markovian multi-armed bandits (MAB) (Gittins et al., 2011). In an MAB, the algorithm is given $n$ Markovian “arms”; at every step the algorithm can pull any one of the arms and receive a reward; the state of the pulled arm then evolves according to the underlying Markov process. Gittins (1979) showed that an index-based policy is optimal for this problem. MAB is usually studied in an infinite horizon setting with discounted or averaged reward. Dumitriu et al. (2003) extended this approach to the finite horizon setting (where Markov chains have terminal states) and Gupta et al. (2019) further extended it to combinatorial selection. Weitzman’s index-policy for Pandora’s Box is a special case of these results.
In a general bandit superprocess, the $n$ arms are Markov Decision Processes (MDPs) comprising multiple possible actions with different costs/rewards. While it is possible to treat the superprocess as a single large MDP, the exponential size of this MDP makes this approach computationally infeasible, and it is desirable to consider algorithms that “solve” each component process separately and combine the solutions together. Whittle (1980) showed that certain bandit superprocesses admit simple index-policies just like the MAB special case. In particular, when each MDP in the superprocess admits an unambiguously optimal static policy, these policies can be composed to obtain a globally optimal policy through appropriately defined indices. However, in the absence of such a strong local optimality condition, the globally optimal policy may take actions that are suboptimal for the MDP they’re taken in.
Much like Whittle’s work, our approach considers a local problem $(\mathcal{M},y)$ for each MDP $\mathcal{M}$ where the algorithm is presented with a deterministic outside option $y$ and at any step can either choose to take an action in $\mathcal{M}$ or accept the outside option $y$. The optimal solution to this local problem generates an optimality curve $f_{\mathcal{M}}(y)$ as a function of the outside option. A primary technical contribution of our work is to convert this optimality curve into a mapping from terminal states of the MDP to “surrogate” costs (Section 3). The surrogate costs essentially allow an algorithm to amortize the costs of actions and pay (a part of) them only when the MDP terminates. We develop a recursive “water filling” algorithm for cost amortization that ensures that good, i.e. low cost or high reward, terminal states are responsible for paying most of the cost share. The overall index of the MDP is then defined to be the cost of the best-case scenario – the lowest possible surrogate cost that is instantiated. In the special case of Markov chains, our water filling amortization exactly recovers the Gittins index.
A key property of the surrogate costs we define is their independence from the actions chosen by the algorithm in the MDP (Lemma 3.3 in Section 3). The corresponding indices can therefore be defined and computed independently of the algorithm’s stochastic trajectory, much like in Gittins or Weitzman’s setting. This allows us to compose these costs and obtain a global bound on the value of the optimal solution. However, the independence comes at a loss – the surrogate costs provide a bound that is not exact for general MDPs (in contrast to Whittle’s results, for example).
Our approach is heavily influenced by the work of Kleinberg et al. (2016) who presented an amortization viewpoint for Weitzman’s index for classical Pandora’s Box, as well as for its extension to Markov chains. The amortization viewpoint is also implicit in the work of Singla (2017), who bounds the optimal cost by considering a phantom “free information” world. Our amortized cost shares reflect a similar argument. We emphasize that while bounds on the optimal solution’s value have been established previously for the single-selection CICS in special cases, a general bound was not known previously. Our bound applies to any feasibility constraint over any bandit superprocess where the constituent MDPs have a DAG-like structure.
Local and semilocal approximation
Next we consider approximations to the CICS through a class of algorithms called “commitment policies”. A commitment policy for an instance of the CICS specifies a (random) action for the algorithm to take at any state of any component MDP. [^5] E.g., in the weighing scale problem, for each $X_{i}$, we may specify targets $t^{i}_{1},t^{i}_{11},t^{i}_{12},$ etc. such that $X_{i}$ is first compared against $t^{i}_{1}$, and then against $t^{i}_{11}$ or $t^{i}_{12}$ depending on whether the first comparison was True or False, and so on and so forth. The commitment for each alternative depends on its own state and not on the states of other processes. This turns each MDP into a (suboptimal) Markov chain. Armed with these commitments, the algorithm faces a combinatorial MAB which, based on the prior discussion, we know how to solve.
Given the characterization of the optimal cost in terms of properties of the local games $(\mathcal{M},y)$ defined above, one reasonable approach is to try to find a commitment policy for each MDP that is approximately optimal for the local game. However, this does not work. Consider, for example, a minimization MDP $\mathcal{M}$ with two actions. The first action leads to a deterministic value of $1$, and the second leads to a stochastic value of 00 or $1$ with probability $1/2$ each. Then, the first action provides a $2$-approximation for the local game $(\mathcal{M},y)$: its expected cost is $\min(y,1)$ whereas the optimal cost of the local game is half of that quantity, achieved by taking the stochastic option. Now consider a single-selection superprocess with $n$ copies of the above MDP. If we commit to the deterministic action in each MDP, our cost is deterministically $1$, whereas taking the stochastic action in each costs us $1/2^{n}$ – an exponential gap!
Scully and Doval (2024) describe how to fix this problem in the context of Pandora’s Box with optional inspection and define a local approximation as follows. Let $f_{\mathcal{M}}(\cdot)$ and $f_{\pi}(\cdot)$ denote the optimality curves of the MDP $\mathcal{M}$ and the commitment policy $\pi$ for $\mathcal{M}$, respectively, in the local game $(\mathcal{M},y)$. The local approximation factor for minimization achieved by $\pi$ for the MDP $\mathcal{M}$ is $\alpha\geq 1$ if we have
$f_{\pi}(\alpha y)\leq\alpha f_{\mathcal{M}}(y)\,\,\forall y\in\real.$ (1)
For maximization, we likewise require $f_{\pi}(\frac{y}{\alpha})\geq\frac{1}{\alpha}f_{\mathcal{M}}(y)$ for all $y$. It then holds that any commitment policy that achieves an $\alpha$-local approximation for each constituent MDP achieves a global $\alpha$-approximation. [^6] Observe that Equation (1) is stronger than simply requiring the commitment policy to be approximately optimal for the local game $(\mathcal{M},y)$ at every possible outside option $y$ (in other words, $f_{\pi}(y)\leq\alpha f_{\mathcal{M}}(y)$). We show in Section 4 that this composition result holds for general CICS.
Informally speaking, local approximation is equivalent to establishing second order stochastic dominance between the surrogate costs of the MDP and the committing policy. We also explore the stronger condition of first order stochastic dominance between surrogate costs, that we call pointwise approximation, which can be easier to establish in some settings. These characterizations allows us to design commitment policies that achieve a good approximation.
In some settings, the local approximation condition can be challenging to satisfy. For example, if $f_{\mathcal{M}}$ is near-linear, then Equation (1) is essentially equivalent to requiring $f_{\mathcal{M}}=f_{\pi}$. [^7] That is, $\pi$ must be unambiguously optimal for $\mathcal{M}$; and no suboptimal policy can provide any local approximation. We therefore explore a weakening of the condition by allowing for additive deviations from (1). We call this property Semilocal Approximation, and show that it leads to improved approximation bounds for some settings (Section 7.3).
Our results
We establish new or improved approximation factors for three variants of the CICS. In each case our results apply to matroid feasibility constraints in the given setting. For the first and second settings, our results extend also to feasibility constraints admitting frugal approximation algorithms as defined in Singla (2017).

### v2

@@ Abstract
We develop a framework for approximate optimization of bandit superprocesses that applies to arbitrary acyclic MDPs with a matroid feasibility constraint. Our framework establishes a bound on the optimal cost through a novel cost amortization; it then couples this bound with a notion of local approximation that allows approximate solutions for each component MDP in the superprocess to be composed without loss into a global approximation.
We use this framework to obtain approximately optimal solutions for several variants of bandit superprocesses for both maximization and minimization. We obtain new approximations for combinatorial versions of the previously studied Pandora’s Box with Optional Inspection and Pandora’s Box with Partial Inspection; the less-studied Additive Pandora’s Box problem; as well as a new problem that we call the Weighing Scale problem.
@@ 1 Introduction
Many real-world settings involve algorithmic decision-making under incomplete information about the underlying input and future outcomes. In some cases, additional information can be acquired at a cost. A fundamental question arises: how should this costly information acquisition be incorporated into the algorithmic process?
Consider, for example, a biotechnology company running multiple drug discovery projects. These projects require upfront investment and entail a series of choices on which directions to pursue before their potential is known and realized. Other examples include an oil company’s search for an optimal drilling site, a construction firm paying for multiple architectural designs before choosing one to pursue, and a manufacturer conducting market research before deciding which products to produce and in what quantity.
A classical model for costly information acquisition is the Pandora’s Box problem, introduced by Weitzman (1979). In this model, an algorithm faces an optimization problem over $n$ stochastic alternatives, each concealed inside a closed box. The algorithm can open a box at a cost to observe the value of the alternative within, and can select an alternative only after its box has been opened. The goal is to minimize the total cost, which includes both the cost of opening the boxes and the values of the chosen alternatives. [^4] For maximization problems, the objective is to maximize the function value minus the cost of opening the boxes. In this section, we will focus our exposition on minimization problems although, as we show later in the paper, our framework and techniques extend to maximization problems as well. The Pandora’s Box model (henceforth, PB) is well understood and admits a simple optimal solution for a broad class of optimization problems. However, it is too simplistic to capture the complexities of real-world information acquisition: it assumes that each alternative can be revealed in only one way and that the algorithm either learns everything or nothing about it—there is no concept of partial information acquisition.
Recent research has explored extensions of PB that allow for multiple modes and stages of information acquisition. [^5] See, for example, Olszewski and Weber (2015); Kleinberg et al. (2016); Singla (2017); Doval (2018); Beyhaghi and Kleinberg (2019); Esfandiari et al. (2019); Gupta et al. (2019); Boodaghians et al. (2020); Aouad et al. (2020); Fu et al. (2022); Beyhaghi and Cai (2022); Berger et al. (2023); Bowers and Waggoner (2024a); Scully and Doval (2024). These generalizations are typically NP-hard (Fu et al., 2022), necessitating the development of approximation algorithms. Beyhaghi and Cai (2024) provide a survey of recent advances in this domain. However, a significant limitation of prior work is that existing solutions are tailored to specific problem variants and do not extend to broader settings. For example, Aouad et al. (2020) address a two-stage inspection process, but their techniques do not generalize to three stages. Similarly, while previous results separately handle cases with partial inspection and optional inspection, it is unclear how to approach an instance that contains alternatives of both kinds.
In this paper, we develop a general framework for modeling arbitrarily complex protocols for information acquisition and designing approximation algorithms for them. Using this framework, we obtain improved approximations for several PB variants and develop algorithms for new models where non-trivial approximations were previously unknown.
A Framework for Modeling Information Acquisition.
We model the process of costly information acquisition for each alternative $i\in[n]$ using a finite-horizon Markov decision process (MDP), denoted by $\mathcal{M}_{i}$. The current state of this MDP encapsulates all information the algorithm has about alternative $i$ and determines the set of available actions for further exploration. Some actions lead to terminal states, where the alternative can be selected. For instance, both a standard optional-inspection PB and a three-stage inspection PB can be modeled as MDPs (see Figure 1). A key assumption in our framework is the independence of these MDPs: exploring one alternative does not affect the state of another.
This framework defines a broad family of optimization problems, which we call Costly Information Combinatorial Selection (CICS). In CICS, we are given $n$ stochastic alternatives, each associated with an MDP $\mathcal{M}_{1},\dots,\mathcal{M}_{n}$. The goal is to select a subset of alternatives satisfying a given feasibility constraint; our work focuses on matroid feasibility constraints. The algorithm iteratively chooses an alternative $i$ and takes an action in $\mathcal{M}_{i}$ to gather information. At any time, it can stop and select a feasible subset of alternatives that have reached terminal states. The objective function is the total expected value of the selected alternatives plus the cost of all actions taken.
Figure 1: Three examples of a Pandora’s Box style information acquisition protocol. (Left:) A classical PB is modeled as a Markov chain with a single costly action (opening the box) resulting in a random terminal state of value $X$. (Middle:) An optional-inspection PB modeled as an MDP; in addition to opening the box, there is a grab action that reveals no information and results in a single terminal state of deterministic value $\operatorname{E}\left[X\right]$. (Right:) A three-stage inspection protocol where the box can be opened directly, or it can be partially explored by sequentially revealing hidden variables $T_{i}$ at a cost.
The primary challenge in solving CICS arises from the interplay between two levels of decision-making: (i) global decision-making, which determines which alternative to explore at each step, and (ii) local decision-making, which dictates the costly actions taken within each alternative’s MDP. These decisions are typically adaptive, responding to the information that was acquired in previous steps. Since CICS generalizes MDPs, we have no hope of obtaining a general purpose solution to the problem. Instead, what we aim for is a decomposition of the problem into smaller pieces that can each be solved independently and composed together into a global solution. We obtain such a decomposition by analyzing a class of algorithms called committing policies.
Committing Policies and the Commitment Gap.
A committing policy for CICS is a policy that, in essence, predetermines all local decisions before making any global decisions. In particular, it preselects an action for each state in every MDP $\mathcal{M}_{i}$, effectively reducing the MDPs to Markov chains. This is valuable because for the class of feasibility constraints we are interested in, the special case of CICS where each MDP $\mathcal{M}_{i}$ is effectively a Markov chain, i.e. has just one action in every state, admits a simple optimal policy (Dumitriu et al., 2003; Gupta et al., 2019). By fixing local decisions in advance, we significantly reduce the complexity of global decision-making. This prompts two questions:
  1. What is the best approximation ratio achievable by committing policies for a given class of CICS problems? We call this ratio the commitment gap.
  2. Can the local decision commitments be made in a composable way? That is, can the commitments for each MDP $\mathcal{M}_{i}$ be determined using only the characteristics of $\mathcal{M}_{i}$, independent of other MDPs?
These questions have been posed and investigated for optional-inspection PB (Beyhaghi and Kleinberg, 2019; Scully and Doval, 2024), but little is known beyond that setting. In this paper, we address both questions by developing a set of local sufficient conditions for bounding the commitment gap in general CICS. By “local”, we mean that our conditions depend only on the properties of each individual MDP $\mathcal{M}_{i}$, without reference to the other MDPs or the feasibility constraint. Furthermore, our conditions yield policies that determine their commitments locally (or, in one case, with a simple global algorithm after local pre-processing).
We introduce two key conceptual contributions that enable bounds on the commitment gap: a method of bounding the optimum, and notions of approximation ratio that can be checked locally for each MDP.
Bounding the Optimum.
Our first conceptual contribution is a lower bound on the cost of the (unrestricted, non-committing) optimum. This bound is analogous to the Whittle integral (Whittle, 1980) for infinite-horizon discounted-reward MDPs. While prior proofs of the Whittle integral exist for the discounted-reward setting (Brown and Smith, 2013; Hadfield-Menell and Russell, 2015), we present an algorithmic proof for the finite-horizon case, which is both simpler and more insightful.
Importantly, our approach extends the notion of surrogate costs from PB literature to general MDPs and connects it to the Whittle integral, offering valuable algorithmic insights into bounding the commitment gap. Surrogate costs were first defined for classical PB by Kleinberg et al. (2016) and Singla (2017), and can be extended easily to arbitrary Markov chains (Weber, 1992; Dumitriu et al., 2003; Gupta et al., 2019). However, the extension to MDPs is highly non-trivial and is a novel contribution of our work. We describe the details in Section 1.1.
Local Approximation.
Our second contribution is the development of a local approximation technique, which expresses the global approximation factor of a committing policy $(\pi_{1},\dots,\pi_{n})$ in terms of the local properties of each constituent policy $\pi_{i}$ relative to its corresponding MDP $\mathcal{M}_{i}$. This concept was first introduced by Scully and Doval (2024) for the optional-inspection PB; we generalize it to arbitrary MDPs.
Crucially, local approximations compose: if each $\pi_{i}$ achieves an $\alpha$-local approximation for $\mathcal{M}_{i}$, then the committing policy $(\pi_{1},\dots,\pi_{n})$ achieves an $\alpha$-global approximation. This observation answers, for example, our aforementioned question of how to compose partial-inspection Pandora’s Boxes with optional-inspection ones within a single instance. Furthermore, our global approximation guarantee extends to any combinatorial setting that admits a “greedy-style” algorithm via techniques similar to (Singla, 2017; Gupta et al., 2019).
Although local approximation can be defined directly with respect to the Whittle integral, as in Scully and Doval (2024), proving guarantees using this definition appears very challenging. One of our main contributions is to connect this concept to surrogate costs in a manner that enables simpler analysis. We develop multiple approaches for bounding local approximation factors that are suitable for different contexts, and demonstrate their use through the applications discussed in Section 1.1.
Example Application: Shortest Paths.
To illustrate the power of our framework, consider the Pandora’s Shortest Path problem introduced by Singla (2017): the edge weights of a given graph correspond to independent Pandora’s boxes and the objective is to accept a set of (open) boxes that form a path between two given vertices $s$ and $t$, while minimizing the total cost of opening boxes plus the cost of the path. As path constraints do not admit frugal algorithms that enable the generalization of Weitzman’s result to combinatorial settings, this problem is notoriously challenging and no approximation results are known.
When the underlying graph is a union of disjoint $s$-$t$ paths, this problem becomes an instantiation of CICS: each path corresponds to an alternative, whose MDP encapsulates the different orders in which the constituent “edge boxes” can be opened. A committing policy, in this case, pre-defines a protocol for the order of inspecting the edges within a path, that can adapt only to the instantiations of the edges from the same path. We show that the commitment gap of this special case of Pandora’s Shortest Paths is at most $2$.
1.1 Our technical contributions and their relationship to prior work
We now describe our technical contributions in more detail in the context of closely related prior work. A thorough discussion of other related work is presented in Section 1.2.
Bandit superprocesses and the Whittle integral
CICS is closely related to a class of sequential decision-making problems called bandit superprocesses (BSPs), introduced by Nash (1973). A BSP is composed from multiple independent MDPs; at every step, the algorithm advances one of the MDPs, receives a (discounted) reward, and the state of the underlying MDP evolves stochastically. A primary difference between the two settings is that a CICS algorithm must terminate by selecting a feasible set of alternatives, whereas a BSP can potentially have an infinite horizon. Moreover, since the states of the MDPs in a CICS model information acquisition, we assume that each constituent MDP is acyclic; our techniques rely on this acyclicity. In contrast, BSPs can be defined over arbitrary MDPs.
In the special case where the MDPs are Markov chains, BSP reduces to the extensively-studied Multi-Armed Bandit (MAB) problem. MAB admits an astonishingly simple optimal solution (Gittins, 1979): every state in each constituent Markov chain is assigned an index independent of the other chains; at every step, the Markov chain with the maximum index in its current state is advanced. Dumitriu et al. (2003) show that this “indexability” extends to the finite-horizon setting as well, and Gupta et al. (2019) further extend it to combinatorial settings. We restate this result with a new proof as Theorem 3.2. Weitzman (1979)’s algorithm for classical PB is essentially a special case of this result.
General BSPs, in contrast, are not indexable (Glazebrook, 1982). [^6] This is because MAB only involves global decision-making and does not require any local decision-making as there is only one way to advance the chosen Markov chain. Nevertheless, the local structure of the problem provides insight into the optimal cost: given an MDP $\mathcal{M}$, Whittle (1980) considers a local problem $(\mathcal{M},y)$, where at every step the algorithm can either advance $\mathcal{M}$ or terminate with a cost of $y$. The values of these local games, formalized as “optimality curves” $f_{\mathcal{M}}(y)$, one for each constituent MDP in the BSP, can be combined to obtain the Whittle integral, a lower bound on the global optimum (Whittle, 1980; Brown and Smith, 2013).
Our work provides a new interpretation for Whittle’s optimality curves by connecting them with a mapping from trajectories in the MDP to “surrogate” costs. The surrogate costs essentially allow an algorithm to amortize the costs of actions and pay (a part of) them only when the MDP terminates. We develop a recursive water filling algorithm for cost amortization that ensures that good, i.e. low cost or high reward, terminal states are responsible for paying most of the cost share. A crucial property we achieve is the independence of the distribution of surrogate costs from the actions chosen by an algorithm in the MDP in the local or global game. In particular, the costs capture the local structure of the MDP without limiting in any way how an algorithm “solves” the MDP. This independence allows us to compose them into a global bound for the optimal (adaptive, non-committing) policy.
Our approach is heavily influenced by Kleinberg et al. (2016) and Singla (2017)’s amortization viewpoint on Weitzman’s index for classical PB, as well as its extension to Markov chains. However, we emphasize that the extension to MDPs while ensuring action independence is an important and novel contribution of our work that enables our approximation bounds.
Commitment gap and local approximation
We follow the approach of Scully and Doval (2024) to establish a bound on the commitment gap by quantifying the performance of each commitment $\pi_{i}$ relative to the optimal policy for MDP $\mathcal{M}_{i}$ in the local game $(\mathcal{M}_{i},y)$. Such an approach helpfully disentangles global and local decision making.
Local guarantees for $(\mathcal{M}_{i},y)$ need to be established with care, however. Whittle (1980) showed that if every $\mathcal{M}_{i}$ admits a commitment $\pi_{i}$ that is optimal for the local game $(\mathcal{M}_{i},y)$ regardless of the value $y$, then the tuple $\mathcal{P}=(\pi_{1},\dotsc,\pi_{n})$ is globally optimal. In other words, unambiguous local optimality composes into global optimality. However, the reverse doesn’t hold: the globally optimal policy might take actions that are unambiguously suboptimal in the local game! (See Appendix C.) Likewise, simply relating the optimal cost $f_{\mathcal{M}_{i}}(y)$ to the cost of the commitment $f_{\mathcal{M}_{i}^{\pi_{i}}}(y)$ at every $y$ is not sufficient, as we demonstrate in Appendix C.
Following Scully and Doval (2024) we say that $\pi_{i}$ $\alpha$-locally approximates $\mathcal{M}_{i}$ if for all $y$ we have $f_{\mathcal{M}_{i}^{\pi_{i}}}(\alpha y)\leq\alpha\cdot f_{\mathcal{M}_{i}}(y)$. Local approximations defined in this manner compose into global guarantees on the commitment gap. We use this approach to obtain new approximations for PB with partial inspections (defined below).
Unfortunately, however, local approximations can be very challenging to prove for general MDPs. We develop a new (but weaker) local condition based on surrogate costs that we call pointwise approximation. This allows us to analyze more complicated MDPs with multiple rounds of actions for which optimality curves are tricky to establish.
Finally, the commitment gap can sometimes be strictly smaller than what can be established through local approximation. We investigate this issue in the context of PB with optional inspection and show that improved results can be obtained through a semi-local argument.
Applications of our framework
We instantiate our framework to obtain new bounds on the commitment gap of four PB extensions. The first three variants are new problems studied in the minimization setting, and we obtain the following results: